Nina borrowed money on 31 August and agreed to pay back the loan on 2 November of the same year. If the discount rate is 18% per year and she received R5 000 on 31 August, what is the value of the loan that Nina has to pay the bank on 2 November?

[1] R5 155,34

[2] R4 844,66

[3] R5 160,32

[4] R5 000,00

[5] None of the above


1
Expert's answer
2021-06-14T16:19:11-0400

Interest on loan="principal\times rate\times time"


"=5000\times \frac{18}{100}\times \frac{30+31+2}{365}"


"=5000\times0.18\times\frac{63}{365}"


"=155.34"


Value of loan=Principal+Interest


"=5000+155.34= R5,155.34"